OpenStudy (1dfanatic):

Which ordered pair is the solution of the system? 6x - 6y = 2 3x + 9y = -7

OpenStudy (adryin):

First try to cancel out x

OpenStudy (adryin):

Elimination can be used here

OpenStudy (adryin):

Multiply the second equation by 2

OpenStudy (1dfanatic):

if i cancel out x wouldnt the second equation turn into 3x?

OpenStudy (adryin):

Yes, which is why you multiply by 2

OpenStudy (adryin):

6x-6y=2 2(3x+9y=-7)

OpenStudy (adryin):

6x-6y=2 6x+18y=-14

OpenStudy (1dfanatic):

im so confused..,